In a simple firing testing, 1.03 g of an explosive material, which has a power index of 211.71

and can produce a gas product up to 9.2 dm3 per gram, was packed in a handgun ammunition

cartridge and triggered to detonate to propel a projectile weighed 2 g out of the barrel. The

efficiency of the energy transferred from the explosion to the projectile is 13 %. Calculate the

propelled bullet velocity in m/s if the heat evolved and gas volume produced in a detonation of

per gram picric acid is 3150 J/g and 0.831 dm3


.


1
Expert's answer
2021-04-29T07:25:27-0400

Energy produced per gram of picric acid, "Q_{picric}=3150\space J/g"

Volume of picric acid, "V_{picric}=0.831\space dm^3"

Power Index "=211.71"

Let energy produced per gram of explosive material be "Q"

Volume of explosive material per gram "=9.2\space dm^3/g"

Volume of explosive material, "(V)" for 1.03 g "=9.2\times1.03=9.476\space dm^3"

Power Index "=\dfrac{QV}{Q_{picric}V_{picric}}"


"Q=\dfrac{211.71\times3150\times0.831}{9.476}"

"Q=58482.765\space J/g"

Energy of Explosion "=58482.765\times1.03=60237.247\space J"

Efficiency of energy transferred from the explosion to the projectile "=13\%"

Energy of projectile "=0.13\times60237.247=7830.84\space J"

Mass of projectile, "m=2\space g=2\times10^{-3}\space kg"

Let velocity of projectile be "v"

"\dfrac{1}{2}mv^2=7830.84"

"v=\sqrt{\dfrac{2\times7830.84}{2\times10^{-3}}}"

"v=2798.36\space m/s"